OpenStudy (anonymous):

simplify the following expression. ( 2 to the third power ) to the fifth power / 2 to the fifteenth power ( / - Divided by ) A. 2 B. 2 to the third power C. 2 to the negative seventh power D. 1

OpenStudy (superflytyguy):

(2^3)^5 / 2^15 2^3 = 8 and 2^15 = 32768 8^5 = 32768 (crazy right?) 32768 / 32768 = 1

OpenStudy (anonymous):

i understood none of that but thanks

OpenStudy (midhun.madhu1987):